Linear ststems
solve using substitution
y=3-2x
y=5+2x

Answer by tutorcecilia(2152)

Substitute one y-equation for the other. In other words, set the equations equal to each other
5+2x=3-2x
Simplify and solve for the x-value
5-3=-2x-2x
2=-4x
-2/4=-4x/4
-2/4=x
Reduce the fractio to lowest terms
-1/2=x
Check by plugging -1/2 back into the equals